OpenStudy (anonymous):

factor 36r-25b^2. I just get 0 on the graphing calculator.

OpenStudy (anonymous):

\[(6\sqrt{r}-5b)(6\sqrt{r}+5b)\] not positive this is the answer that you should arrive at

OpenStudy (anonymous):

\[\left( 6\sqrt{r}+5b \right)\left( 6\sqrt{r} -5b\right)\]